OpenStudy (anonymous):

2/3y+1/4=3/2y+3/4 y=?

OpenStudy (anonymous):

3/20

OpenStudy (anonymous):

2/3 y - 3/2y = 3/4 - 1/4 =4/6y - 9/6 y = 2/4 -5/6 y = 2/4 y = 2/4 * (-6/5) y = -3/5 Maybe I am doing something wrong since I didn't get 3/20

OpenStudy (anonymous):

which is right

OpenStudy (anonymous):

I took denominators as (3y+(1/4)) & (2y+(3/4)) that is why i got 3/20

OpenStudy (anonymous):

i made it 3/5

OpenStudy (anonymous):

If not swirlingsara is correct

OpenStudy (anonymous):

i took them same as swirling sara

OpenStudy (anonymous):

k then whats correct answer

OpenStudy (anonymous):

it was -3/5

OpenStudy (anonymous):

is that coorrect dogface? denominators are y and 4?

OpenStudy (anonymous):

yes -3/5

OpenStudy (anonymous):

k